# UniBEV: Multi-modal 3D Object Detection with Uniform BEV Encoders for Robustness against Missing Sensor Modalities

### v2

@@ Abstract
Multi-sensor object detection is an active research topic in automated driving, but the robustness of such detection models against missing sensor input (modality missing), e.g., due to a sudden sensor failure, is a critical problem which remains under-studied. In this work, we propose UniBEV, an end-to-end multi-modal 3D object detection framework designed for robustness against missing modalities: UniBEV can operate on LiDAR plus camera input, but also on LiDAR-only or camera-only input without retraining. To facilitate its detector head to handle different input combinations, UniBEV aims to create well-aligned Bird’s Eye View (BEV) feature maps from each available modality. Unlike prior BEV-based multi-modal detection methods, all sensor modalities follow a uniform approach to resample features from the native sensor coordinate systems to the BEV features. We furthermore investigate the robustness of various fusion strategies w.r.t. missing modalities: the commonly used feature concatenation, but also channel-wise averaging, and a generalization to weighted averaging termed Channel Normalized Weights. To validate its effectiveness, we compare UniBEV to state-of-the-art BEVFusion and MetaBEV on nuScenes over all sensor input combinations. In this setting, UniBEV achieves $52.5\%$ mAP on average over all input combinations, significantly improving over the baselines ($43.5\%$ mAP on average for BEVFusion, $48.7\%$ mAP on average for MetaBEV). An ablation study shows the robustness benefits of fusing by weighted averaging over regular concatenation, and of sharing queries between the BEV encoders of each modality. Our code will be released upon paper acceptance.
@@ I INTRODUCTION
The perception system of an intelligent vehicle typically relies on multiple sensors[1, 2], including LiDARs and cameras, to take advantage of their individual strengths and complementary nature for robust object detection. For example, cameras provide rich texture information while LiDAR delivers dense point clouds with accurate geometric information. Most works on multi-sensor models focus only on optimal detection performance when all sensors are available. However, ideally a model could also be used when the input of one of its sensors is missing, i.e., modality missing, without any retraining. A unified model to handle both multi-sensor and single-sensor inputs would facilitate gracefully degradation of its perception system in case of catastrophic sensor failure (e.g., broken connector), but also provide flexibility by supporting diverse hardware configurations (e.g., vehicles with different sensors). This work therefore focuses on the design of a ‘robust’ multi-sensor object detection model, which in this context refers to the trained model’s ability to fuse camera and LiDAR information for object detection, but also to operate on only a single modality.
@@
Figure 1: Comparison of our UniBEV with other relevant works. (a). BEVFusion [3] fuses multi-modal BEV features extracted from two separate branches with concatenation. (b) MetaBEV [4] fuses multi-modal BEV features extracted from two separate branches with a fusion module consisting of several deformable attention layers. (c) Our UniBEV extracts multi-modal BEV features from their original coordinate systems with uniform BEV encoders and fuses the BEV features with the CNW module. C and L in the figure represent the input from cameras and LiDAR, respectively. Q denotes the unified queries for BEV encoders.
@@
Recent state-of-the-art (SotA) in multi-sensor object detection for self-driving exploits dense Bird’s-Eye view (BEV) features as an intermediate representation to integrate multi-sensor information, which can then be used by a generic object detection head. BEVFusion [5] pioneers fusing multi-modal BEV features for LiDAR and cameras. It employs two separate branches to independently extract BEV features from each modality, and subsequently fuses these features through concatenation, as shown in Fig. 1 (a). Notably, the designs of the camera and LiDAR branches in BEVFusion are non-uniform: The camera branch relies on a (type of) Lift-Splat-Shoot (LSS) [6] component to explicitly predict the depth distribution of the image features, and map them from their native camera coordinates to the spatial BEV coordinates. In contrast, the LiDAR branch already natively expresses its features in spatial coordinates and thus does not apply additional transformations to encode its BEV features. This may lead to the misalignment of camera and LiDAR BEV features as they are extracted in distinct ways. Recently MetaBEV [4] improves over BEVFusion by replacing its concatenation by a learnable module with several deformable attention layers to better align the features, while still keeping BEVFusion’s BEV feature encoder approach, see Fig. 1 (b). Since the feature misalignment is not really solved, the gain of the fusion inference is modest.
@@
  • We propose UniBEV, a multi-modal 3D object detection model aiming for robustness against modality missing. Its detector head works on fused BEV feature maps from all modalities. Unlike prior detectors that fuse multi-modal BEV-feature maps, UniBEV uses a uniform approach across all modalities to encode the sensor-specific feature maps into shared BEV feature space, which facilitates the alignment between modalities. Experiments on nuScenes show UniBEV is more robust than SotA BEVFusion and MetaBEV.
@@
  • We investigate the impact of shared BEV queries between the BEV encoders of all modalities, compared to using separate queries, and demonstrate consistent small improvements for all input combinations.
@@ II RELATED WORK
Every sensor type has specific limitations for real-world driving scenarios, hence multi-modal 3D object detection has gained great attention in recent years, especially the fusion between cameras and LiDAR. Nevertheless, the alignment between camera and LiDAR features is challenging as they are defined in different coordinates. Some methods fuse multi-modal features directly from their native coordinates with some specially designed components, such as attention [7]. [8, 9, 10, 11] DeepFusion [12] is the representative work among them, it simply performs cross-attention on multi-modal features with LiDAR features as queries and with camera features as keys and values. FUTR3D [13] brings the idea of DETR3D [14] and Object DGCNN [15] into the multi-modal domain. With the prior knowledge of camera extrinsics, FUTR3D leverages the shared object queries interacting with multi-view image features and LiDAR BEV features to sample instance-level features.
Other SotA methods build a unified intermediate BEV representation to align and fuse multi-modal features.  [16, 17] A benefit of such BEV representations is they can serve various tasks through distinct network heads, such as simultaneous object detection and BEV map segmentation[5, 18, 19]. BEVFusion [5, 3] deploys Lift-Splat-Shoot (LSS) [6] to predict image depth distributions and project the image features in BEV, and deploys a regular point voxelization method, such as PointPillars [20] or CenterPoint [21], to extract BEV features from LiDAR point clouds. The multi-modal BEV feature maps are fused by concatenation. MetaBEV [4] upgrades the fusion module of BEVFusion to a deformable attention-based fusion block. [22] uses a simple summation fusion module to integrate cross-modal BEV features and achieve convincing performance at a long distance. In this study, our primary emphasis is on the alignment within the BEV feature domain to address the challenges posed by missing sensor scenarios.
@@ III-A Feature Extractors
For the initial feature extraction, UniBEV has a similar design to previous works, relying on common image/point cloud backbones[5, 3, 13, 24]. Following previous works [18, 14], images from the $V$ camera views are fed into an image backbone, such as ResNet-101 [25]. resulting in image features $F^{i}_{\mathbf{C}}\in\mathbb{R}^{H_{\mathbf{C}}\times W_{\mathbf{C}}\times N_{\mathbf{C}}}$ for $1\leq i\leq V$, where $H_{\mathbf{C}}\times W_{\mathbf{C}}$ is the resolution of the feature map in the native image coordinates, and $N_{\mathbf{C}}$ is the feature dimension. Similarly, the LiDAR scan is processed by a regular point cloud backbone, such as VoxelNet [26], which voxelizes and extracts grid-shaped features in Bird’s Eye view $F_{\mathbf{L}}\in\mathbb{R}^{H_{\mathbf{L}}\times W_{\mathbf{L}}\times N_{\mathbf{L}}}$, where $H_{\mathbf{L}}$, $W_{\mathbf{L}}$, and $N_{\mathbf{L}}$ are the spatial shape and feature dimensions of the features.
@@ III-B Uniform BEV Feature Encoders
After feature extraction, $F_{\mathbf{L}}$ and $F_{\mathbf{C}}$ are still represented in different coordinate systems. $F_{\mathbf{L}}$ is expressed in 3D spatial coordinates similar to the target BEV space, while $F_{\mathbf{C}}$ uses 2D image coordinates. Existing methods generally further transfer the image features into the Bird’s Eye view with LSS [6] and simply fuse the two BEV features through concatenation. [5, 27] We argue that the difference in network architecture between these branches may affect the alignment of the camera and LiDAR BEV features. Furthermore, concatenating features requires zero-filling when one modality is missing. As a result, the decoder head would operate on BEV features that are highly different depending on the available inputs, which may impact its robustness to a missing modality. UniBEV therefore implements a uniform design for all sensor modalities for better aligned BEV features, as explained next.
@@
Projection: The BEV spatial locations $R$ are projected to the native spatial coordinate system of each modality’s feature map, as shown in Fig. 2, similar to FUTR3D [13]  [^1] Recall FUTR3D is not a BEV-based detector, nor does it use deformable attention to sample camera features.. Namely, for the feature map of each camera $i=\{1,2,...,V\}$, the 3D points $R$ are projected to its 2D image-based coordinates $R^{i}_{\mathbf{C}}=\mathsf{P}_{\mathbf{C}}(R,P^{i})$ using the known camera extrinsics expressed by its homogeneous projection matrix $P^{i}$. Similarly, $R_{\mathbf{L}}=\mathsf{P}_{\mathbf{L}}(R)$ projects the references to the LiDAR feature map’s spatial coordinates, for instance to scale the spatial resolution, though in practice often $\mathsf{P}_{\mathbf{L}}$ is an identity function.
Encoding: Finally, each modality’s BEV feature map is constructed using 3 layers of deformable self-attention and deformable cross-attention between the BEV queries and sensor feature maps. The feature map at the first layer of the camera BEV encoder, $F^{BEV^{\prime}}_{\mathbf{C}}$, is obtained by summing over the all views where a reference is visible, and over all $D$ locations for each query [18],
@@ III-C Fusion Module: Channel Normalized Weights
In the majority of multi-modal 3D object detection approaches [5, 3, 27, 24, 13], concatenation along the feature dimension is utilized as the fusion method to combine features from different modalities, with the aim of preserving a maximum amount of information. However, if one considers scenarios where a sensor input is missing, concatenation fusion becomes ill-defined as the number of fused channels provided to the decoder should not change. One must compensate the absent input by filling the fused BEV features with placeholder values, typically all zeros.
@@
We also propose a generalization of average fusion which we call Channel Normalized Weights (CNW). CNW learns a $N$-dimensional weight vector $A_{m}$ for each modality $m$ which remains fixed after training. The $i$-th element $A_{m}(i)$ indicates the relative importance of modality $m$ for the fused result of the $i$-th channel. Before fusion, weights $A_{m}$ are normalized (denoted $\overline{A}_{m}$) over all available sensor modalities such that they sum to 1 per channel. Thus with two modalities, LiDAR and camera, $\overline{A}_{\mathbf{C}}(i),\overline{A}_{\mathbf{L}}(i)=softmax(A_{\mathbf{C}}(i),A_{\mathbf{L}}(i))$, s.t.
@@
It is easy to see that CNW reduces to average fusion when all the learned channel weights in $\overline{A}_{\mathbf{C}}$ and $\overline{A}_{\mathbf{L}}$ approach $\frac{1}{2}$. On the other hand, CNW can also reflect concatenation fusion by allowing channels in the fused output to only take information from one modality, if those channels’ learned weights approach 00 or $1$ only. Intuitively, CNW adds a small number of learnable parameters to give the model more flexibility between these special cases, allowing it to optimize the relative importance of each modality for fusion, and still allowing meaningful values for the single modal input. Our experimental results shall show UniBEV constructs BEV features with a similar magnitude distribution for each modality, ensuring that our CNW discerns the importance of different channels rather than a random scale function.
@@ III-D Detection Head and Modality Dropout Strategy
Following previous works [18, 14, 30, 13], we cast the bounding box detection as a set prediction problem and adopt the decoder of BEVFormer [18] for 3D object detection task. To train the model for sensor missing failure, we deploy the common Modality Dropout (MD) training strategy [23, 24, 4]. Thus during training we drop with a probability $p_{md}$ the BEV features of one random modality, $F^{BEV}_{\mathbf{C}}$ or $F^{BEV}_{\mathbf{L}}$.
TABLE I: Evaluation results on the nuScenes val set (best/ second best). Columns indicate the test input modalities: L+C = LiDAR and cameras, L = only LiDAR, C = only cameras. Non-fusion models provided for completeness, using ”-” where they do not apply.
@@ IV-A Implementation Details
Dataset and Metrics. We train and evaluate our approach on the nuScenes dataset [2]. nuScences is a large-scale multi-modal driving dataset, which includes 6 cameras and a 32-beam LiDAR in the sensor suite. As we target robustness against missing input modalities, we report the test performance on LiDAR+camera, on LiDAR-only and on camera-only, and report common mean Average Precision (mAP) and nuScenes detection score (NDS) [2]. As a summary metric for ‘robustness to modality missing’, we also report for both metrics the average performance over all these three possible sensor inputs. For example, our summary mAP is simply,
@@
Model. We use ResNet-101 [25] with FPN [31] as UniBEV’s camera feature extractor, and VoxelNet [26] as its LiDAR feature extractor. CNW is the default fusion approach. The grid shape of the unified query is set to $200\times 200$ with $N=256$.
Baselines. Our first multi-sensor baseline is BEVFusion [5], and uses their implementation which includes a more powerful image backbone, Dual-Swin-Tiny [32], and CenterPoint[21] head. Since BEVFusion uses concatenation for fusion, we use zero-filling for MD. Our second fusion baseline is the recent MetaBEV [4], which improves the concatenation approach of BEVFusion to a deformable attention-based fusion module. Our MetaBEV implementation uses the same backbones and detector head as UniBEV [^2] MetaBEV is currently arXiv only, no code has been released yet. We reproduced their BEV-encoder and fusion strategies based on the paper. Their multi-task learning strategy was not implemented for fair comparison..
@@
Training Details. Our model is trained end-to-end manner for 36 epochs. For fair comparison, we do not adopt the CBGS [33] augmentation for our model nor for the reproduced baselines. The Modality Dropout (MD) probability for all models is $p_{md}=0.5$. The image and point cloud backbones are initialized with the weights of FCOS3D [34] and CenterPoint[21], respectively. Our model is implemented in the open-sourced MMDetection3D[35].

### v3

@@ Abstract
Multi-sensor object detection is an active research topic in automated driving, where the robustness of such detection models against missing sensor input (modality missing), e.g., due to a sudden sensor failure, is a critical problem that remains under studied. In this work, we propose UniBEV, an end-to-end multi-modal 3D object detection framework designed for robustness against missing modalities: UniBEV can operate on LiDAR plus camera input, but also on LiDAR-only or camera-only input without retraining. To facilitate its detector head to handle different input combinations, UniBEV aims to create well-aligned Bird’s Eye View (BEV) feature maps from each available modality. Unlike prior BEV-based multi-modal detection methods, all sensor modalities follow a uniform approach to resample features from the original sensor coordinate systems to the BEV features. We furthermore investigate the robustness of various fusion strategies w.r.t. missing modalities: the commonly used feature concatenation, but also channel-wise averaging, and a generalization to weighted averaging termed Channel Normalized Weights. To validate its effectiveness, we compare UniBEV to state-of-the-art BEVFusion and MetaBEV on nuScenes over all sensor input combinations. In this setting, UniBEV achieves better performance than these baselines for all input combinations. An ablation study shows the robustness benefits of fusing by weighted averaging over regular concatenation, and of sharing queries between the BEV encoders of each modality. Our code is available at https://github.com/tudelft-iv/UniBEV.
@@ I INTRODUCTION
The perception system of an intelligent vehicle typically relies on multiple sensors[1, 2], including LiDARs and cameras, to take advantage of their individual strengths and complementary nature for robust object detection. For example, cameras provide rich texture information while LiDAR delivers dense point clouds with accurate geometric information. Most works on multi-sensor models focus only on optimal detection performance when all sensors are available. However, ideally a model could also be used when the input of one of its sensors is missing, i.e., modality missing, without any retraining. A unified model to handle both multi-sensor and single-sensor inputs would facilitate gracefully degradation of its perception system in case of catastrophic sensor failure (e.g., broken connector) without needing to load a new set of model parameters, but also provide flexibility by supporting diverse hardware configurations (e.g., vehicles with different sensors). This work therefore focuses on the design of a ‘robust’ multi-sensor object detection model, which in this context refers to the trained model’s ability to fuse camera and LiDAR information for object detection, but also to operate on only a single modality.
@@
Figure 1: Comparison of our UniBEV with other relevant works. (a). BEVFusion [3] fuses multi-modal BEV features extracted from two separate branches with concatenation. (b) MetaBEV [4] fuses multi-modal BEV features extracted from two separate branches with a fusion module consisting of several deformable attention layers. (c) Our UniBEV extracts multi-modal BEV features from their original coordinate systems with uniform BEV encoders and fuses the BEV features with the CNW module. C and L in the figure represent the input from cameras and LiDAR.
@@
Recent state-of-the-art (SotA) in multi-sensor object detection for self-driving exploits dense Bird’s-Eye view (BEV) features as an intermediate representation to integrate multi-sensor information, which can then be used by a generic object detection head. BEVFusion [5] pioneers fusing multi-modal BEV features for LiDAR and cameras. It employs two separate branches to independently extract BEV features from each modality, and subsequently fuses these features through concatenation, as shown in Fig. 1 (a). Notably, the designs of the camera and LiDAR branches in BEVFusion are non-uniform: The camera branch relies on a (type of) Lift-Splat-Shoot (LSS) [6] component to explicitly predict the depth distribution of the image features, and map them from their camera coordinates to the spatial BEV coordinates. In contrast, the LiDAR branch already natively expresses its features in spatial coordinates and thus does not apply additional transformations to encode its BEV features. This may lead to the misalignment of camera and LiDAR BEV features as they are extracted in distinct ways. Recently MetaBEV [4] improves over BEVFusion by replacing its concatenation with a learnable module with several deformable attention layers to better align the features, while still keeping BEVFusion’s BEV feature encoder approach, see Fig. 1 (b). Since the feature misalignment is not really solved, the gain of the fusion inference is modest.
@@
  • We propose UniBEV, a multi-modal 3D object detector designed for robustness against missing modalities. It follows a uniform approach across all modalities to encode the sensor-specific features into a shared BEV feature space to facilitate alignment between modalities. UniBEV exhibits a more robust performance than its baselines against modality missing failure without needing to load a different set of model parameters.
  • When taking multi-modal data as input, UniBEV outperforms SotA multi-modal methods on the nuScenes dataset. For fair benchmarking, we reimplemented the closed-source MetaBEV and evaluated all multi-modal baselines under similar training conditions (e.g. same hardware, no data augmentation).
@@ II RELATED WORK
Every sensor type has specific limitations for real-world driving scenarios, hence multi-modal 3D object detection has gained great attention in recent years, especially the fusion between cameras and LiDAR. Nevertheless, the alignment between camera and LiDAR features is challenging as they are defined in different coordinates. Some methods fuse multi-modal features directly from their native coordinates with some specially designed components, such as attention [7, 8, 9, 10, 11]. DeepFusion [12] is the representative work among them, it simply performs cross-attention on multi-modal features with LiDAR features as queries and with camera features as keys and values. FUTR3D [13] brings the idea of DETR3D [14] and Object DGCNN [15] into the multi-modal domain. With the prior knowledge of camera extrinsics, FUTR3D leverages the shared object queries interacting with multi-view image features and LiDAR BEV features to sample instance-level features.
Other SotA methods build a unified intermediate BEV representation to align and fuse multi-modal features [16, 17]. A benefit of such BEV representations is they can serve various tasks through distinct network heads, such as simultaneous object detection and BEV map segmentation [5, 18, 19, 20, 6]. BEVFusion [5, 3] deploys Lift-Splat-Shoot (LSS) [6] to predict image depth distributions and project the image features in BEV, and deploys a regular point voxelization method, such as PointPillars [21] or CenterPoint [22], to extract BEV features from LiDAR point clouds. The multi-modal BEV feature maps are fused by concatenation. MetaBEV [4] upgrades the fusion module of BEVFusion to a deformable attention-based fusion block. [23] uses a simple summation fusion module to integrate cross-modal BEV features and achieve convincing performance at a long distance. In this study, our primary emphasis is on the alignment within the BEV feature domain to address the challenges posed by missing sensor scenarios.
@@ III-A Feature Extractors
For the initial feature extraction, UniBEV has a similar design to previous works [18, 14], relying on common image/point cloud backbones[5, 3, 13, 25]. Images from $V$ camera views are fed into an image backbone, such as ResNet [26], resulting in image features $F^{i}_{\mathbf{C}}\in\mathbb{R}^{H_{\mathbf{C}}\times W_{\mathbf{C}}\times N_{\mathbf{C}}}$ for $1\leq i\leq V$, where $H_{\mathbf{C}}\times W_{\mathbf{C}}$ is the resolution of the feature map in the native image coordinates, and $N_{\mathbf{C}}$ is the feature dimension. Similarly, the LiDAR scan is processed by a regular point cloud backbone, such as VoxelNet [27], which voxelizes and extracts grid-shaped features in Bird’s Eye view $F_{\mathbf{L}}\in\mathbb{R}^{H_{\mathbf{L}}\times W_{\mathbf{L}}\times N_{\mathbf{L}}}$, where $H_{\mathbf{L}}$, $W_{\mathbf{L}}$, and $N_{\mathbf{L}}$ are the spatial shape and feature dimensions of the features.
@@ III-B Uniform BEV Feature Encoders
After feature extraction, $F_{\mathbf{L}}$ and $F_{\mathbf{C}}$ are still represented in different coordinate systems. $F_{\mathbf{L}}$ is expressed in 3D spatial coordinates similar to the target BEV space, while $F_{\mathbf{C}}$ uses 2D image coordinates. Existing methods generally further transfer the image features into the Bird’s Eye view with LSS [6] and simply fuse the two BEV features through concatenation [5, 28]. We argue that the difference in network architecture between these branches may affect the alignment of the camera and LiDAR BEV features. Furthermore, concatenating features requires zero-filling when one modality is missing. As a result, the decoder head would operate on BEV features that are highly different depending on the available inputs, which may impact its robustness to a missing modality. UniBEV therefore implements a uniform design for all sensor modalities for better aligned BEV features, as explained next.
@@
Projection: The BEV spatial locations $R$ are projected to the original spatial coordinate system of each modality’s feature map, as shown in Fig. 2, similar to FUTR3D [13]  [^1] Recall FUTR3D is not a BEV-based detector, nor does it use deformable attention to sample camera features.. Namely, for the feature map of each camera $i=\{1,2,...,V\}$, the 3D points $R$ are projected to its 2D image-based coordinates $R^{i}_{\mathbf{C}}=\mathsf{P}_{\mathbf{C}}(R,P^{i})$ using the known camera extrinsics expressed by its homogeneous projection matrix $P^{i}$. Similarly, $R_{\mathbf{L}}=\mathsf{P}_{\mathbf{L}}(R)$ projects the references to the LiDAR feature map’s spatial coordinates, for instance to scale the spatial resolution, though in practice often $\mathsf{P}_{\mathbf{L}}$ is an identity function.
Encoding: Finally, each modality’s BEV feature map is constructed using 3 layers of deformable self-attention and deformable cross-attention between the BEV queries and sensor feature maps. The feature map at the first layer of the camera BEV encoder, $F^{BEV^{\prime}}_{\mathbf{C}}$, is obtained by summing over all views where a reference is visible, and over all $D$ locations for each query [18],
@@ III-C Fusion Module: Channel Normalized Weights
In the majority of multi-modal 3D object detection approaches [5, 3, 28, 25, 13], concatenation along the feature dimension is utilized as the fusion method to combine features from different modalities, to preserve a maximum amount of information. However, if one considers scenarios where a sensor input is missing, concatenation fusion must compensate for the absent input to ensure the number of channels provided to the decoder does not change. For instance, by filling the fused BEV features with placeholder values, typically all zeros.
@@
We also propose a generalization of average fusion which we call Channel Normalized Weights (CNW). The key idea is that the reliability of each channel in the feature map may differ from sensor to sensor, and we should account for this when sensor measurements can be fused. CNW therefore has as learnable parameters a $N$-dimensional vector $A_{m}$ for each modality $m$, which remains fixed after training. The $i$-th element $A_{m}(i)$ indicates the relative importance of modality $m$ for the fused result of the $i$-th channel. Before fusion, weights $A_{m}$ are normalized (denoted $\overline{A}_{m}$) over all available sensor modalities such that they sum to 1 per channel. Thus with two modalities, LiDAR and camera, $\overline{A}_{\mathbf{C}}(i),\overline{A}_{\mathbf{L}}(i)=softmax(A_{\mathbf{C}}(i),A_{\mathbf{L}}(i))$, s.t.
@@
It is easy to see that CNW reduces to average fusion when all the learned channel weights in $\overline{A}_{\mathbf{C}}$ and $\overline{A}_{\mathbf{L}}$ approach $\frac{1}{2}$. On the other hand, CNW can also reflect concatenation fusion by allowing channels in the fused output to only take information from one modality if those channels’ learned weights approach 00 or $1$ only. Intuitively, CNW adds a small number of learnable parameters to give the model more flexibility between these special cases, allowing it to optimize the relative importance of each modality for fusion, and still allowing meaningful values for the single modal input. Our experimental results shall show UniBEV constructs BEV features with a similar magnitude distribution for each modality, ensuring that our CNW discerns the importance of different channels rather than a random scale function.
@@ III-D Detection Head and Modality Dropout Strategy
Following previous works [18, 14, 31, 13], we cast the bounding box detection as a set prediction problem and adopt the decoder of BEVFormer [18] for 3D object detection task. To train the model for sensor missing failure, we deploy the common Modality Dropout (MD) training strategy [24, 25, 4]. Thus during training we drop with a probability $p_{md}$ the BEV features of one of the modalities, either $F^{BEV}_{\mathbf{C}}$ or $F^{BEV}_{\mathbf{L}}$. Furthermore, in case we do drop one of the modalities, $p_{L}$ indicates the probability of keeping LiDAR, while $p_{C}=1-p_{L}$ is the probability of keeping the cameras. Thus, the overall probability of keeping both sensors is $1-p_{md}$, of only LiDAR is $p_{md}\cdot p_{L}$, and of only cameras is $p_{md}\cdot p_{C}=p_{md}\cdot(1-p_{L})$.
TABLE I: Evaluation results on the nuScenes val set (best/ second best). Columns indicate the test input modalities: L+C = LiDAR and cameras, L = only LiDAR, and C = only cameras. Non-fusion models are provided for completeness, using ”-” where they do not apply. Note: we trained all models ourselves, to ensure equal training strategies, and since [4] has no public code. This leads to lower performance compared to reported benchmark results[3, 4], especially due to a lack of data augmentation.
@@ IV-A Implementation Details
Dataset and Metrics. We train and evaluate our approach on the nuScenes dataset [2]. nuScences is a large-scale multi-modal driving dataset, which includes 6 cameras and a 32-beam LiDAR in the sensor suite. As we target robustness against missing input modalities, we report the test performance on LiDAR+camera, on LiDAR-only, and on camera-only, and report common mean Average Precision (mAP) and nuScenes detection score (NDS) [2]. As a summary metric for ‘robustness to modality missing’, we also report for both metrics the average performance over all these three possible sensor inputs. For example, our summary mAP is simply,
@@
Model. We use ResNet-101 [26] with FPN [32] as UniBEV’s camera feature extractor and VoxelNet [27] as its LiDAR feature extractor. CNW is the default fusion approach. The grid shape of the unified query is set to $200\times 200$ with $N=256$.
Baselines. Our first multi-sensor baseline is BEVFusion [5], and uses their implementation which includes a more powerful image backbone, Dual-Swin-Tiny [33], and CenterPoint[22] head. Since BEVFusion uses concatenation for fusion, we use zero-filling for MD. Our second fusion baseline is the recent MetaBEV [4], which improves the concatenation approach of BEVFusion to a deformable attention-based fusion module. Our MetaBEV implementation uses the same backbones and detector head as UniBEV  [^2] The code of MetaBEV has not been released yet. We reproduced their BEV-encoder and fusion strategies based on the paper. Their multi-task learning strategy was not implemented for a fair comparison..
@@
Training Details. Our model is trained in an end-to-end manner for 36 epochs. Due to the utilization of various tricks in baseline methods and the absence of publicly available code, accurately reproducing the performance reported in their papers proves to be challenging. For a fair comparison, all the baselines are retrained with the same data pipeline as our model without data augmentation techniques, such as CBGS [34]. Every baseline model underwent training on four Nvidia A40 GPUs, with the entire training duration for each model spanning roughly one week.
Unless stated otherwise, for all models the MD probability is $p_{md}=0.5$, and the probability for keeping each modality is identical, i.e., $p_{L}=p_{C}=0.5$. Therefore, in the whole training process, on average $50\%$ of iterations are trained with multi-modal inputs, $25\%$ with LiDAR-only inputs, and $25\%$ with camera-only inputs.
The image and point cloud backbones are initialized with the weights of FCOS3D [35] and CenterPoint[22], respectively. Our model is implemented in the open-sourced MMDetection3D[36].